Assume that a,b are real numbers. Show a^2 +b^2=0 implies a=b=0
1
Expert's answer
2010-09-24T06:51:24-0400
a² >=0, b^2>=0, a^2 +b^2 = 0 => a^2 = 0, b^2 = 0 => a = b